# ControlAR: Controllable Image Generation with Autoregressive Models

### v1

@@ 1 Introdution
 [^1] Equal contribution. Works was done during Zongming Li’s internship at vivo AI Lab. [^2] Corresponding author (xgwang@hust.edu.cn).
Recent advancements in image generation have led to the emergence of various models that leverage text-to-image diffusion models  (Saharia et al., 2022; Ho et al., 2022; Rombach et al., 2022; Podell et al., 2023) to generate high-quality visual content. Among them, several works (Zhang et al., 2023a; Li et al., 2024b; Qin et al., 2023b) such as ControlNet (Zhang et al., 2023a), have explored adding conditional controls to text-to-image diffusion models and allowed for image generation according to the precise spatial controls, e.g., edges, depth maps, or segmentation masks. The control-to-image diffusion models have impressively enhanced the versatility of these models in applications ranging from creative design to augmented reality.
@@ 2.3 Controllable Image Generation
Relying solely on textual prompts is insufficient for conveying distinctive artistic style or precise detail during T2I image generation. Some methods (Gal et al., 2022; Ruiz et al., 2023; Zhang et al., 2023b) attempt to capture concepts from example images that are not easily described through text to guide image generation, a task known as personalization for controllable generation. Represented by ControlNet (Zhang et al., 2023a) and T2I-Adapter (Mou et al., 2024), work in this area utilizes the spatial structure of the image, such as edges, segmentation masks, depth maps, etc., to enable spatial control in the generation process. Subsequently, UniControl (Qin et al., 2023b), Uni-ControlNet (Zhao et al., 2024), and ControlNet++ (Li et al., 2024b) further extend this realm, focusing on condition encoder design and optimization of training strategies. Furthermore, GlueGen (Qin et al., 2023a) pairs a multi-modal encoder with a stable diffusion model for sound-to-image generation. Controllable generation based on autoregressive image generation models has been less explored. ControlVAR (Li et al., 2024c) employs next-scale prediction to jointly model control and image, but is still different from next-token prediction in autoregressive generation. Our objective is to fully harness the capabilities of autoregressive models and explore a general and efficient paradigm for controllable image generation using autoregressive models.
@@ 3.2 Unified Conditional Decoding
Autoregressive image generation models leverage the two-phase generation process, including the prefilling and decoding (Pope et al., 2023; Kwon et al., 2023), where prefilling processes the prompt tokens (or control tokens) and stores them in the KV Cache (Pope et al., 2023), and then decoding follows next-token prediction and aims to generate the output tokens (e.g., image tokens). In ControlAR, we bring the condition into the decoding phase by adding the control condition token to the image token, which we refer to as conditional decoding. Specifically, modify Eq. 2 as follows:
$p(\mathbf{q})=\prod_{t=1}^{h\cdot w}p(q_{t}|q_{1}+C_{2},q_{2}+C_{3},\ldots,q_{t-1}+C_{t},c+C_{1}),$ (4)
where $C$ represents the control condition token generated from the control image. The number of control condition tokens is the same as the number of image tokens expected to be generated. It is worth noting that we use displacement by one position when adding the control condition tokens to the sequence, which allows the model to make autoregressive predictions with control information corresponding to the next image token.
@@ 3.3 Controllable Autoregressive Model
We propose a lightweight control encoder to transform control image to control condition tokens. In contrast to previous approaches such as ControlNet (Zhang et al., 2023a) and T2I-Adapter (Mou et al., 2024), we utilize the Vision Transformer (ViT) (Dosovitskiy, 2020) for feature extraction of control images. We believe that a ViT model, pre-trained on a large amount of data, is more adept at modeling sequences than a randomly initialized CNN network. For the class-to-image task on ImageNet, we use ViT-S to initialize our control encoder. Additionally, for the text-to-image task, we employ DINOv2-S (Oquab et al., 2023) as the initialization scheme for the control encoder. Further details on this are available in section 4.3. Notably, our ControlAR achieves efficient controllable generation with a control encoder comprising only about 22M parameters, which is significantly less than the 361M parameter count of ControlNet++.
@@ 4.2 Experimental Results
Instead of uniformly resizing the controls and images to 512 $\times$ 512, we adopt a set of resolutions ranging from 384 to 1024 and keep the image sequence length to no more than 2304 to train our Multi-Resolution ControlAR. Direct end-to-end controllable generation using MR-ControlAR preserves the detailed features of the control image and avoids the loss of information due to scaling. We show the difference between these two approaches in Fig. 6 (a), and perform hed edge control generation experiments on images with different resolution ratios in the validation set of MultiGen-20M. Experimental results show that after multi-resolution training, MR-ControlAR can ensure that the generation of images with different resolution ratios is not impaired. We show the visualization at different resolutions in Fig. 1.
@@ 4.3 Ablation Studies
In Tab. 4, we conduct experiments using different encoders (or pre-training schemes) towards different controls, including canny edge, depth map, and hed edge. Firstly, we follow T2I-Adapter (Mou et al., 2024) and design a vanilla convolutional control encoder with 4 consecutive residual blocks (He et al., 2016) and a total downsample ratio of 16. The vanilla CNN-based control encoder contains 21.8M parameters, which has similar parameters with ViT-S (Dosovitskiy, 2020). Further, we explore the effects of using pre-trained ViTs as our control encoder and adopt ViTs with different pre-trained schemes, i.e., the ImageNet-supervised (Dosovitskiy, 2020) and self-supervised (Oquab et al., 2023). For our experiments, we employ LlamaGen-B’s C2I model to conduct experiments on the ImageNet (Deng et al., 2009) for canny edge and depth map conditions, while using LlamaGen-XL’s T2I model to carry out experiments on the MultiGen-20M (Qin et al., 2023b) for the hed edge condition. As depicted in the table, different control encoders exhibit varying performance across different datasets, with the ViT-S demonstrating superior suitability for the ImageNet dataset, and the DINOv2-S yielding better results on the MultiGen-20M dataset. Furthermore, we evaluate the performance of larger encoders such as DINOv2-B, and the outcomes reveal that higher parameter counts enable our method to achieve superior results.
@@ A.1 Implementation Details
A.2 Discussion
@@
A.3 More Visualizations
More visualization results under different conditions of control are shown in Fig. 7 8 9 10 11. We alse show some visualization comparison of ControlAR and MR-ControlAR at different resolution in Fig. 12 and Fig. 13.

### v3

@@ 1 Introdution
Recent advancements in image generation have led to the emergence of various models that leverage text-to-image diffusion models  (Saharia et al., 2022; Ho et al., 2022; Rombach et al., 2022; Podell et al., 2023) to generate high-quality visual content. Among them, several works (Zhang et al., 2023a; Li et al., 2024b; Qin et al., 2023b) such as ControlNet (Zhang et al., 2023a), have explored adding conditional controls to text-to-image diffusion models and allowed for image generation according to the precise spatial controls, e.g., edges, depth maps, or segmentation masks. The control-to-image diffusion models have impressively enhanced the versatility of these models in applications ranging from creative design to augmented reality. In addition, these methods are also helpful for more difficult video generation tasks, such as ControlVideo (Zhang et al., 2023b).
@@ 2.3 Controllable Image Generation
Relying solely on textual prompts is insufficient for conveying distinctive artistic style or precise detail during T2I image generation. Some methods (Gal et al., 2022; Ruiz et al., 2023; Zhang et al., 2023c) attempt to capture concepts from example images that are not easily described through text to guide image generation, a task known as personalization for controllable generation. Represented by ControlNet (Zhang et al., 2023a) and T2I-Adapter (Mou et al., 2024), work in this area utilizes the spatial structure of the image, such as edges, segmentation masks, depth maps, etc., to enable spatial control in the generation process. Subsequently, UniControl (Qin et al., 2023b), Uni-ControlNet (Zhao et al., 2024), and ControlNet++ (Li et al., 2024b) further extend this realm, focusing on condition encoder design and optimization of training strategies. Furthermore, GlueGen (Qin et al., 2023a) pairs a multi-modal encoder with a stable diffusion model for sound-to-image generation. Controllable generation based on autoregressive image generation models has been less explored. ControlVAR (Li et al., 2024c) employs next-scale prediction to jointly model control and image, but is still different from next-token prediction in autoregressive generation. Our objective is to fully harness the capabilities of autoregressive models and explore a general and efficient paradigm for controllable image generation using autoregressive models.
@@ 3.2 Unified Conditional Decoding
Autoregressive image generation models leverage the two-phase generation process, including the prefilling and decoding (Pope et al., 2023; Kwon et al., 2023), where prefilling processes the prompt tokens (or control tokens) and stores them in the KV Cache (Pope et al., 2023), and then decoding follows next-token prediction and aims to generate the output tokens (e.g., image tokens). In ControlAR, we bring the condition into the decoding phase by adding the control condition token to the image token, which we refer to as conditional decoding. Specifically, we describe it as follows:
$S_{out}=\mathcal{F}(S_{in}+\mathcal{P}(C))=\mathcal{F}([c+C_{1},I_{1}+C_{2},I_{2}+C_{3},...,I_{i-1}+C_{i}]),$ (4)
where $\mathcal{F}$ represents a single sequence layer modeling process in the generative network, $\mathcal{P}$ is the projection function, $S_{in}$ and $S_{out}$ are the input sequence and output sequence of each layer respectively, c is the class or text token, $I_{i}$ is the image token, and $C$ is the control condition sequence. It is worth noting that we use displacement by one position when adding the control condition tokens to the sequence, which allows the model to make autoregressive predictions with control information corresponding to the next image token.
@@ 3.3 Controllable Autoregressive Model
We propose a lightweight control encoder to transform control image to control condition tokens. In contrast to previous approaches such as ControlNet (Zhang et al., 2023a) and T2I-Adapter (Mou et al., 2024), we utilize the Vision Transformer (ViT) (Dosovitskiy, 2020) for feature extraction of control images. We believe that a ViT model, pre-trained on a large amount of data, is more adept at modeling sequences than a randomly initialized CNN network. For the class-to-image task on ImageNet, we use ViT-S to initialize our control encoder. Additionally, for the text-to-image task, we employ DINOv2-S (Oquab et al., 2023) as the initialization scheme for the control encoder. Further details on this are available in section 4.3. Notably, our ControlAR achieves efficient controllable generation with a control encoder comprising only about 22M parameters, resulting in an additional computational effort of only about 0.05T MACs for 512×512 resolution.
@@ 4.2 Experimental Results
Instead of uniformly resizing the controls and images to 512 $\times$ 512, we adopt a set of resolutions to train our Multi-Resolution ControlAR. Specifically, we randomly sample the height and width of the training data from 384 to 1024 with a minimum interval of 16, and the image can be resized when it satisfies (H/16)×(W/16$\leq$2304. In addition, we need to adjust the parameter settings of the rotational position encoding in the generative network by simply increasing its maximum sequence length to 2304. Direct end-to-end controllable generation using MR-ControlAR preserves the detailed features of the control image and avoids the loss of information due to scaling. We show the difference between these two approaches in Fig. 6 (a), and perform hed edge control generation experiments on images with different resolution ratios in the validation set of MultiGen-20M. Experimental results show that after multi-resolution training, MR-ControlAR can ensure that the generation of images with different resolution ratios is not impaired. We show the visualization at different resolutions in Fig. 1.
@@ 4.3 Ablation Studies
In Tab. 4, we conduct experiments using different encoders (or pre-training schemes) towards different controls, including canny edge, depth map, and hed edge. Firstly, we follow T2I-Adapter (Mou et al., 2024) and design a vanilla convolutional control encoder with 4 consecutive residual blocks (He et al., 2016) and a total downsample ratio of 16. The vanilla CNN-based control encoder contains 21.8M parameters, which has similar parameters with ViT-S (Dosovitskiy, 2020). Further, we explore the effects of using pre-trained ViTs as our control encoder and adopt ViTs with different pre-trained schemes, i.e., the ImageNet-supervised (Dosovitskiy, 2020) and self-supervised (Oquab et al., 2023). For our experiments, we employ LlamaGen-B’s C2I model to conduct experiments on the ImageNet (Deng et al., 2009) for canny edge and depth map conditions, while using LlamaGen-XL’s T2I model to carry out experiments on the MultiGen-20M (Qin et al., 2023b) for the hed edge condition. As depicted in the table, different control encoders exhibit varying performance across different datasets. The reason for this phenomenon is the different pre-training data for the two models. ViT-s is obtained by pre-training on ImageNet and thus is more advantageous for C2I tasks that are also trained on ImageNet. DINOv2-s, on the other hand, is pre-trained on a larger and more diverse data such as LVD-142M, and thus will be more suitable for T2I tasks trained on MultiGen20M, which is also a diverse text-image paired dataset. Furthermore, we evaluate the performance of larger encoders such as DINOv2-B, and the outcomes reveal that higher parameter counts enable our method to achieve superior results.
@@ A.1 Implementation Details
A.2 More experimental explorations
Comparison with recent work.
We have added some quantitative comparative results with recent work including OmniGen (Xiao et al., 2024) and Lumina-mGPT (Liu et al., 2024), as shown in the Tab. 9. The results for segmentation task are measured on the validation set of ADE20K (Zhou et al., 2017), and the results for canny, hed and depth are measured on the validation set of MultiGen-20M (Qin et al., 2023b). OmniGen uses iterative denoising diffusion for image generation, while lumina-mGPT uses autoregressive prediction. Although Lumina-mGPT has a much larger number of parameters than our ControlAR, it does not perform particularly well on the controllable generation task. Our ControlAR provides a good solution for autoregressive controllable image generation and our method does not require any adjustments to the structure of the generative network or modifications to the length of the sequences, which means that we can easily migrate our ControlAR to other autoregressive image generation models, such as Lumina-mGPT.
Table 9: Quantitative comparison with recent works.
[…]
OmniGen 3.8B 44.23 35.54 82.37 28.54
Lumina-mGPT 7B 25.02 29.99 78.21 55.25
Ours 0.8B 39.95 37.08 85.63 29.01
Adjustable control strength.
Given the diversity of image structures, we sometimes do not want the spatial structure of the generated image to be identical to the input control. To achieve this, it is only necessary to skip the operation of fusing the control condition token with the image token with a probability of 50% when training ControlAR. Such an approach ensures ControlAR’s generative capability in the absence of control image inputs. At the same time, multiplying the control condition token by a control strength factor $\alpha$ during inference changes the degree of control of the generated result. When $\alpha$ is 1, ControlAR will generate an image exclusively based on the control condition, while when $\alpha$ is 0, the generated results will be related only to the text prompt. Fig. 7 shows the visualizations using edges as the control image and adjusting the control strength.
Figure 7: Visualization with different control strength factor $\alpha$.
Arbitrary-Resolution Generation Without Condition Image.
We conduct a more in-depth exploratory study on resolution control in the absence of specific condition image. We can generate a grayscale map of the corresponding resolution according to the desired height and width, this grayscale map consists of a number of 16 × 16 small squares, and the grayscale value of each row decreases from left to right, the left most 255, the right most 0. This grayscale image is the condition image that determines the resolution. Thanks to the strong positional dependence of the control decoding strategy between the image token and the control condition token, the model only needs to generate a sequence as long as the control condition sequence. And since the grayscale value of each row is decreasing from left to right, the model can easily know when it is necessary to switch to the next row. We have verified the feasibility of this approach on a small experimental scale. We show some visualization results in Fig. 8. Using resolution-aware prompts to control the resolution as in Lumina-mGPT requires the constant generation of $<end-of-line>$ tokens during the prediction of the image and the eventual prediction of $<end-of-image>$ token. This approach requires the model to make its own decisions about where to make line breaks and where to end generation, but our ControlAR is directly telling the model where to make line breaks and end generation. We only need to fine-tune the weights based on LlamaGen-XL (512×512) on about 1M text-image paired data for 30k steps to achieve a good arbitrary resolution generation capability without specific control image. This proves that our ControlAR can be a very effective strategy for controlling resolution.
A.3 Discussion
Limitation.
Figure 8: Arbitrary-Resolution generation without condition image. The grayscale map on the left is the condition image generated according to the desired resolution.
Failure Cases.
ControlAR performs well on the conditional consistency of controllable generation of spatial structures. But because of this, the generated images are sometimes less controlled by the text prompt, especially when the textual prompt conflicts with the spatial structure of the control image. We use depth-to-image and canny-to-image as examples in Fig. 9. When there is a large difference between the text prompt and the original image, it might fail to generate images according to the text prompt. In ControlAR, we can use the control factor to adjust the strength of spatial control, thereby aligning the generated results with the text and mitigating this conflict. However, the conflict between text prompts and spatial controls is a common issue in current control-to-image generation models, including ControlNet (Zhang et al., 2023a) and ControlNet++ (Li et al., 2024b). As shown in Fig. 9, neither ControlNet nor ControlNet++ can generate images that strictly follow the text prompts. Moreover, ControlNet++ introduces additional supervision to facilitate alignment between the generated image and spatial controls, which weakens the influence of the text prompt as shown in the case of canny-to-image. This phenomenon reflects that there may be some confrontation between the structural freedom of the generated image and the conditional consistency. In order to improve the diversity of generated images, we believe that we need to explore suitable training strategies to achieve the effect of being able to adjust the intensity of control during the inference phase, and to resolve the possible contradiction between text alignment and conditional consistency, which are important directions in future research.
Figure 9: Failure cases of current ControlAR. When the text prompt conflicts with the control image, the generated result tends to ignore the text prompt. Adjusting the control strength factor $\alpha$ can alleviate this problem.
A.4 Discussion
@@
A.5 More Visualizations
More visualization results under different conditions of control are shown in Fig. 10 11 12 13 14. We alse show some visualization comparison of ControlAR and MR-ControlAR at different resolution in Fig. 15 and Fig. 16.